# Machine Learning in Fundamental Physics Natural Image Classification via Quasi-Cyclic Graph Ensembles and Random-Bond Ising Models at the Nishimori Temperature

### v1

@@ 7 Conclusions
We introduced a physics-inspired embedding that maps CNN features onto spins of a Random-Bond Ising Model defined on MET-QC-LDPC graphs. Trapping sets create Ihara–Bass zeta poles that appear as isolated eigenvalues of the non-backtracking operator, encoding $\mathbb{Z}/2$-torsion, Betti numbers and bordism obstructions. A Nishimori-temperature estimator evaluates the Bethe–Hessian eigenvalue at three temperatures, fits a quadratic interpolant and applies one Newton step, reaching convergence in $\approx 9$ Arnoldi iterations—six times faster than bisection. Graph design combines permanent and Bethe-permanent bounds with ACE/EMD to suppress trapping sets, yielding sparse spherical/toridal graphs. These compress MobileNetV2’s 1280-dimensional features to 32 dimensions for ImageNet-10 and 64 for ImageNet-100, achieving top-1 accuracies of 98.7% and 82.7% while cutting parameters and memory by factors of 40 and 20. The proposed method unites statistical mechanics, coding theory and topology into a sparse QC-LDPC embedding that exceeds baselines with lower cost, enabling efficient DNN.
Appendix A Trapping sets as topological defects in the spectral embedding
@@
In this appendix, we give a completely rigorous step-by-step proof that any trapping-set subgraph $\mathcal{T}$ of a QCLDPC Tanner graph creates a low-energy (negative) mode of the Bethe–Hessian $H_{\beta,J}$ and, equivalently, produces a pole of the Ihara–Bass zeta function. The argument is split into four formal statements.
@@
  • $H\in\{0,1\}^{m\times n}$ – binary incidence (parity-check) matrix of a QC-LDPC code; rows = check nodes, columns = variable nodes.
@@
$B_{ij}^{\mathcal{T}}(\beta)$ $=\frac{\tanh(\beta J_{ij})}{1-\tanh^{2}(\beta J_{ij})},$ $i\neq j,\;(i,j)\in E_{\mathcal{T}},\$4pt]D^{\mathcal{T}}_{ii}(\beta)$ $=1+\sum_{j:(i,j)\in E_{\mathcal{T}}}\frac{\tanh^{2}(\beta J_{ij})}{1-\tanh^{2}(\beta J_{ij})}.$
@@
Let $C\subseteq\mathcal{T}$ be a simple cycle of length $\ell\geq 2$ satisfying (6) on each of its edges. Then the restriction $H_{\beta_{N},J}^{C}$ possesses exactly one negative eigenvalue; all other eigenvalues are non-negative. The corresponding (up to a gauge transformation) eigenvector is $v=(1,-1,1,-1,\dots)\in\mathbb{R}^{\ell},$ i.e. it alternates sign along the cycle.
Proof.
Because the Tanner graph of an LDPC code is bipartite, every cycle $C$ is even; write $\ell=2r$. Hence a gauge transformation $x_{i}\;\mapsto\;s_{i}\,x_{i},\qquad s_{i}\in\{\pm 1\},$ can be chosen so that after the change of variables every edge of $C$ has positive coupling: $\tanh(\beta J_{ij})=+\sqrt{1-\varepsilon},\qquad\varepsilon:=1-\tanh^{2}(\beta J_{ij})>0.$ The parameter $\varepsilon$ depends on the edge, but for all edges of $C$ we have $\varepsilon\downarrow 0$ as $\beta\to\beta_{N}$. For an edge $(i,j)\in C$ we compute $\frac{\tanh^{2}(\beta J_{ij})}{1-\tanh^{2}(\beta J_{ij})}=\frac{1-\varepsilon}{\varepsilon},\qquad\frac{\tanh(\beta J_{ij})}{1-\tanh^{2}(\beta J_{ij})}=\frac{\sqrt{1-\varepsilon}}{\varepsilon}.$ All other edges (those not belonging to $C$) contribute bounded terms that remain finite as $\varepsilon\to 0$; we denote their total contribution by a matrix $R_{\varepsilon}$ satisfying $R_{\varepsilon}=O(1)$. Let $A_{C}$ be the adjacency matrix of the cycle (entries equal to 1 on the two neighboring positions, zero otherwise). Using Lemma A.1 we obtain for the restriction to $C$: $H_{\beta,J}^{C}=\Bigl{(}1+\frac{2(1-\varepsilon)}{\varepsilon}\Bigr{)}I-\frac{\sqrt{1-\varepsilon}}{\varepsilon}A_{C}+R_{\varepsilon}.$ Factorizing $1/\varepsilon$ and expanding $\sqrt{1-\varepsilon}=1+O(\varepsilon)$ gives the asymptotic expansion
@@
Because $\mu_{0}=0$, 8 gives $\lambda_{0}(\varepsilon)\to-1$, whereas for every $k\neq 0$ we have $\mu_{k}>0$ and thus $\lambda_{k}(\varepsilon)\to+\infty$ as $\varepsilon\downarrow 0$. Consequently, in the limit $\beta\to\beta_{N}$ there is exactly one negative eigenvalue (tending to 1-1) and all remaining eigenvalues are non-negative. The eigenvector belonging to $\mu_{0}=0$ is the constant vector in the gauge- transformed basis. Undoing the gauge transformation multiplies each component by the sign factor $s_{i}$, which yields precisely the alternating pattern $v=(1,-1,1,-1,\dots)$. Since the cycle length is even, this vector is well defined. ∎
@@
Proof.
Decompose $\mathcal{T}$ into its $c(\mathcal{T})$ connected components $\mathcal{T}^{(1)},\dots,\mathcal{T}^{(c)}$. Because $H_{\beta_{N},J}^{\mathcal{T}}$ is block-diagonal with respect to this decomposition, the total number of negative eigenvalues is the sum over the components. Fix a component $\mathcal{C}$. Choose a spanning forest $F$ of $\mathcal{C}$; the edges of $F$ form a tree and thus contribute no negative eigenvalue (its Bethe–Hessian reduces to a positive semidefinite matrix, see Lemma A.1). Every edge $e\in E(\mathcal{C})\setminus F$ creates exactly one independent cycle when added to $F$. By Proposition 1 each such cycle produces one negative eigenvalue of $H_{\beta_{N},J}^{\mathcal{C}}$, and different edges generate linearly independent cycles, hence distinct negative eigenvalues. Therefore, $q_{\mathcal{C}}=|E(\mathcal{C})|-|V(\mathcal{C})|+1.$ Summing over all components gives $q_{\mathcal{T}}=\sum_{c=1}^{c(\mathcal{T})}\!\bigl{(}|E(\mathcal{T}^{(c)})|-|V(\mathcal{T}^{(c)})|+1\bigr{)}=|E(\mathcal{T})|-|V(\mathcal{T})|+c(\mathcal{T})=\beta_{1}(\mathcal{T}),$ which is precisely the claimed equality. ∎
@@
We recall the non-backtracking (Hashimoto) matrix and the Ihara–Bass zeta function.
@@
For a finite graph $\Gamma$ let $\mathcal{B}(\Gamma)$ be its non-backtracking operator. The Ihara–Bass zeta function is $\zeta_{\Gamma}(u)=\prod_{[C]}\bigl{(}1-u^{\,\ell(C)}\bigr{)}^{-1},$ where the product runs over all primitive (non-repeating) cycles $[C]$ and $\ell(C)$ denotes their length.
A classical identity (see e.g. [7]) links $\zeta_{\Gamma}$, the non-backtracking matrix, and the Bethe–Hessian:
@@
Corollary A.2.1.
Let $\Gamma$ be any graph. If for some $u_{c}$ we have $\det(H_{\beta,J(u_{c})})=0$, then $u_{c}$ is a pole of the Ihara–Bass zeta function. In particular, every negative eigenvalue identified in Theorem A.2 produces a pole at $u_{c}=\tanh(\beta_{N}J_{e})$, where $J_{e}$ is the coupling on the corresponding cycle edge.
Proof.
Equation (9) shows that $\zeta_{\Gamma}(u)$ diverges exactly when $\det(H_{\beta,J(u)})=0$. By Theorem A.2, at the Nishimori temperature each independent cycle forces an eigenvalue of $H_{\beta_{N},J}$ to cross zero (from negative for $\beta>\beta_{N}$ to positive for $\beta<\beta_{N}$). Hence the corresponding $u_{c}=\tanh(\beta_{N}J)$ is a pole of $\zeta_{\Gamma}(u)$. ∎
The exact correspondence proved above relies on the strict Nishimori condition $\tanh^{2}(\beta_{N}J_{ij})=1$ for every edge belonging to a cycle. In practice this holds only for uniform couplings $|J_{ij}|=\infty$ or when the distribution of $J$ is chosen so that the Nishimori line coincides with $\tanh^{2}=1$. Nevertheless, the mechanism is robust: Even if $\tanh^{2}(\beta J_{ij})<1$ but close to 1 on a cycle, the two terms in (5) become large with opposite sign. The matrix $H_{\beta,J}^{C}$ then has one very small eigenvalue, which becomes negative as soon as $\beta$ exceeds a critical value that is close to $\beta_{N}$. Consequently, away from the exact Nishimori line the number of negative eigenvalues need not equal $\beta_{1}(\mathcal{T})$, but it is still bounded below by the number of balanced even cycles and above by the total number of independent cycles. In other words, $\beta_{1}(\mathcal{T})-\delta\leq q_{\mathcal{T}}\leq\beta_{1}(\mathcal{T})$ for a small defect $\delta$ that vanishes as the Nishimori condition is approached. The Ihara–Bass identity (9) holds for any graph, irrespective of the values of $J$. Therefore trapping sets always generate poles of $\zeta_{\Gamma}$; at the exact Nishimori temperature these poles correspond to eigenvalue crossings of the Bethe–Hessian. Thus the result proved in Steps 1–4 is a sharp statement for the idealized Nishimori point, while the qualitative picture (cycles $\rightarrow$ low-energy modes $\rightarrow$ zeta-function poles) persists under realistic, approximate conditions.

### v2

@@ 6 Conclusions
We introduced a physics‑inspired embedding that maps high‑dimensional CNN descriptors onto spins of a Random‑Bond Ising Model defined on specially constructed MET‑QC‑LDPC graphs. The central theoretical insight is a spectral–topological correspondence: trapping sets in the Tanner graph generate poles of the Ihara–Bass zeta function, which appear as isolated eigenvalues of both the non‑backtracking operator and the Bethe–Hessian. These spectral signatures encode $\mathbb{Z}/2$‑torsion, Betti numbers and bordism obstructions—topological defects in the feature manifold. On the algorithmic side we proposed a quadratic‑Newton estimator for the Nishimori temperature. By evaluating the smallest eigenvalue of the Bethe–Hessian at three trial temperatures, fitting a quadratic interpolant and applying a single Newton correction, convergence is reached in roughly nine Arnoldi iterations, yielding about a sixfold speed‑up over conventional bisection while preserving numerical accuracy. The graph design pipeline combines permanent and Bethe‑permanent bounds with ACE/EMD optimisation to suppress low‑genus trapping sets, producing two families of sparse graphs (spherical and toroidal). These structures compress the 12801280‑dimensional MobileNetV2 features to $32$ dimensions for a ten‑class ImageNet subset and to $64$ dimensions for a hundred‑class subset. Despite this compression we obtain top‑1 accuracies of 98.7% on ImageNet‑10 and 84.9% on ImageNet‑100, while reducing FLOPs by up to 29-times. Thus, by unifying statistical‑mechanics, coding‑theoretic graph design, and algebraic‑topological analysis, the proposed method delivers compact yet accurate classifiers that are well suited for energy‑constrained edge deployment.
Приложение A Topological Signatures of Trapping Sets
$x_{0}$$x_{1}$$x_{2}$$x_{3}$$c_{0}$$c_{1}$$c_{2}$$c_{3}$$c_{4}$
$x_{0}$$x_{1}$$x_{2}$$x_{3}$$c_{0}$$c_{1}$$c_{2}$$c_{3}$$c_{4}$$c_{5}$$c_{6}$$c_{7}$$c_{8}$$c_{9}$$c_{10}$$c_{11}$$c_{12}$
Рис. 8: Graphical representation of TS(4, 2) (left), TS(4, 6) (center) and TS(9, 2) (right)
A.1 Spectral manifestation of trapping sets
[…]
A.2 Ihara–Bass zeta function
For a finite graph $G$ let $\mathcal{B}$ denote the non‑backtracking operator. The Ihara–Bass zeta function is defined by $\zeta_{G}(u)=\prod_{[C]}\bigl(1-u^{\,\ell(C)}\bigr)^{-1},$ where the product runs over equivalence classes $[C]$ of primitive cycles and $\ell(C)$ is the length of a cycle, [26] . Poles of $\zeta_{G}$ occur at reciprocals of eigenvalues of $\mathcal{B}$; therefore each closed cycle contributes a factor $(1-u^{\ell})^{-1}$. When $G$ contains a trapping set, the associated cycles generate poles in $\zeta_{G}(u)$, which manifest as isolated eigenvalues of both $\mathcal{B}$ and the Bethe–Hessian $H_{\beta,J}$. These poles can be interpreted as $\mathbb{Z}/2$‑torsion elements: they correspond to non‑trivial 11‑cycles ($\beta_{1}>0$) in the underlying feature manifold. Removing trapping sets eliminates the corresponding poles and reduces the number of low‑energy modes.
A.3 Bordism triviality and topological obstructions
Let $G=(V,E)$ be a simple graph and let $\phi:V\rightarrow\mathbb{R}^{d}$ embed its vertices as points of a smooth manifold $M_{G}:=\phi(V)\subset\mathbb{R}^{d}$, [27]. We call $G$ bordism‑trivial if $M_{G}$ is null–bordant in the appropriate bordism group, i.e. there exists a compact (d+1)(d+1)‑manifold $W$ with $\partial W=M_{G}\qquad(\text{or }\partial W=M_{G}\sqcup(-M_{G})\text{ for the unoriented case}),$ and all characteristic numbers of $M_{G}$ vanish: $w_{k}(M_{G})=0,\;k>0,\qquad p_{j}(M_{G})=0,\;j\geq 1.$ Equivalently $M_{G}]=0$ in the oriented bordism group $\Omega^{SO}_{d}$ (and in $\Omega^{O}_{d}$ when orientation is ignored). For a multi‑class problem let $\{M_{i}\}_{i=1}^{K}$ be the manifolds obtained from the feature vectors of each class. A bordism obstruction between classes $i$ and $j$ occurs if the disjoint union $M_{ij}:=M_{i}\sqcup(-M_{j})$ is not null‑bordant in $\Omega^{SO}_{d}$ (or $\Omega^{O}_{d}$). A convenient detection criterion is the non‑vanishing of a characteristic number, e.g. $\exists\,k>0:\;w_{k}[M_{ij}]\neq 0\quad\text{or}\quad\exists\,j:\;p_{j}[M_{ij}]\neq 0.$ If such an obstruction exists for any pair of classes, no graph‑based embedding can achieve perfect separation in the high‑dimensional feature space, [28].
A.4 Higher signatures and KK–theory
Given a cell complex $\mathcal{M}_{G}$ derived from the graph, let $L_{k}=d_{k-1}^{*}d_{k-1}+d_{k}d_{k}^{*}$ be the combinatorial Hodge Laplacian acting on kk‑cochains, [29]. The kernel dimension satisfies $\dim\ker L_{k}=\beta_{k},$ so that spectral information directly yields Betti numbers.
For any cohomology class $x\in H^{*}(B\pi_{1}(\mathcal{M}_{G});\mathbb{Q})$, the higher signature $\sigma_{x}(\mathcal{M}_{G})=\bigl\langle L(\mathcal{M}_{G})\smile x,\,[\mathcal{M}_{G}]\bigr\rangle$ pairs the Hirzebruch LL‑class with the fundamental class, [30]. By the Novikov conjecture, $\sigma_{x}\!\!\pmod{2}$ detects non‑trivial $\mathbb{Z}/2$‑torsion, hence odd‑length trapping cycles render some higher signature non‑zero.
The Dirac operator on the bipartite graph, $\mathcal{D}=\begin{pmatrix}0&A_{\!vn}\\ A_{\!vn}^{\!\top}&0\end{pmatrix}$, has spectrum symmetric about zero, [31, 32]. Its analytic index equals the difference of dimensions of positive and negative eigenspaces, which, by the Atiyah–Singer index theorem, coincides with $\beta_{0}-\beta_{1}$, [33]. Thus a non‑zero index signals the presence of topological defects (e.g. odd cycles).
In KK‑theory, the Kasparov matrix $D_{\mathrm{Kas}}=\begin{pmatrix}0&S&T\\ S^{\!\top}&0&0\\ T^{\!\top}&0&0\end{pmatrix}$, built from suitable incidence matrices $S$ and $T$, satisfies: $K_{0}=\dim\ker D_{\mathrm{Kas}}^{2}$ (counts zero‑modes, i.e. $\beta_{0}$); $K_{1}=\operatorname{rank}D_{\mathrm{Kas}}\bmod 2$ (detects $\mathbb{Z}/2$ torsion, equivalent to the Kervaire invariant), [22] . Non‑trivial KK‑groups therefore provide a concise algebraic signature of the topological obstruction introduced by a trapping set, [34].
The continuous genus is a scalar number that captures how “twisted’’ or “curved’’ a graph $G$ (or, more generally, a network built from points and edges) is, by looking at the spectrum of a matrix derived from the graph $\widehat{A}(H)=\frac{1}{2\sqrt{n_{V}}}\Biggl(\sum_{\lambda_{i}\in\Lambda^{+}}\!\sqrt{\lambda_{i}}-\sum_{\lambda_{j}\in\Lambda^{-}}\!\sqrt{-\lambda_{j}}\Biggr)$, where $n_{V}$ are number of vertices in the graph, [35]. The experimental study of how topological invariants influence spectral embedding is presented in Sec. A.6.
A.5 Permanent and Bethe–permanent bounds
The permanent of an $m\times m$ matrix $\mathbf{B}=[b_{j,i}]_{j,i}$ over a commutative ring is $\operatorname{perm}(\mathbf{B})=\sum_{\sigma\in S_{m}}\;\prod_{j=1}^{m}b_{\,j,\sigma(j)},$ i.e. the determinant without the sign factor $\operatorname{sgn}(\sigma)$. Exact evaluation of (A.5) is #P‑complete, consequently a number of approximation and fast‐evaluation schemes have been proposed [36]. The minimum Hamming distance $d_{\min}$ of a QC‑LDPC code can be bounded by permanents of its weight matrix [37]. For a parity‑check matrix $H(x)$ let $\mathbf{A}=\operatorname{wt}(H(x))$ denote the entrywise Hamming weight. The code distance upper bound, $TS(a=d_{\min},0)$:
where $\operatorname{min}^{*}$ denotes the minimum over all (v+1)(v+1)‑subsets. Replacing the permanent by its Bethe approximation yields an upper bound on the minimum pseudoweight, which controls the harmfulness of $TS(a,b>0)$, [38]. The Bethe‑permanent of a non‑negative matrix $\mathbf{B}$ is defined as $\operatorname{perm}_{\!B}(\mathbf{B})=\exp\!\Bigl(-\min_{q\in\mathcal{Q}}\;F_{\rm Bethe}(q)\Bigr)$, where $F_{\rm Bethe}$ is the Bethe free energy and $\mathcal{Q}$ denotes the set of factorised marginals, [39]. In our construction we maximize weight for both bounds during protograph and parity-check optimization using EMD/ACE maximization, (Alg. 2, [7]), thereby suppressing low‑weight TS and improving the Bethe–Hessian spectrum, Eq. 5.
A.6 Influence of Topological Invariants on Spectral Embedding
For the purpose of illustration we consider six frequently occurring trapping sets $TS(4,2)$, $TS(4,6)$, $TS(9,2)$, $TS(13,6)$, $TS(26,20)$, $TS(28,22)$. Table 3 collects the most relevant topological and spectral quantities for each of them. From a spectral‑embedding perspective a trapping set can be viewed as a locally dense cluster of edges that is attached to only a few low‑degree check nodes. Such a configuration produces negative eigenvalues of the Bethe–Hessian matrix $H_{r}$ (see the column “# neg. $\lambda(H_{1})$”) and, more importantly, gives rise to low‑energy modes that are poorly separated by any embedding based on the graph Laplacian. Consequently, graphs containing many of these substructures generate highly correlated features, which in turn can dramatically deteriorate the performance of downstream classifiers.
Таблица 3: Topological and spectral invariants of Trapping sets.
Invariant $TS(4,2)$ $TS(4,6)$ $TS(26,20)$ $TS(9,2)$ $TS(13,6)$ $TS(28,22)$
$\rho$ (spectral radius) 1.618 4.000 2.7545 8.9168 10.4154 13.5644
$r_{\rm crit}=\sqrt{\rho}$ 1.272 2.000 1.6597 2.9861 3.2273 3.6830
$\{\lambda(H_{1})<0\}$ (negative modes at $r=1$) 1 2 1 1 0 0
$K_{0}$ (zero‑dimensional Betti number) 1 5 7 12 17 45
$K_{1}$ (one‑dimensional Betti number) 1 1 1 0 1 1
Kervaire invariant $\kappa$ 1 1 1 0 0 0
$w_{2}$ (Stiefel–Whitney) 1 1 1 1 1 1
Mod‑2 Betti $(\beta_{0},\beta_{1})$ (1,0) (1,0) (1,0) (1,0) (1,0) (1,0)
Bordism obstruction 0 0 0 0 0 1
To assess the practical impact of these invariants we built two graphs that serve as adjacency matrices in a convolutional architecture trained on the ImageNet‑100 benchmark. Optimized irregular MET QC‑LDPC toric graph. The protograph matrix (3.6 average column weight) is
lifted with a circulant size of $520$, [7]. During the construction we deliberately eliminated harmful trapping sets; the remaining trapping sets is the relatively benign $TS(26,20)$, which possesses a single negative Bethe–Hessian eigenvalue and a modest continuous genus (see Table 3). After a Bayesian classifier we obtained Top‑1 accuracy: $0.6724\pm 2.46\times 10^{-6}$; Top‑3 accuracy: $0.8582\pm 8.53\times 10^{-7}$. Thus, with an average degree of only $3.6$ we achieve performance comparable to much denser near regular graphs. Non-optimized irregular MET QC‑LDPC toric graph (same $\bar{w}=3.6$) was built without filtering out low‑genus, high‑mode traps. Consequently it contains the sets $TS(4,2)$, $TS(9,2)$, $TS(13,6)$ and $TS(28,22)$. These introduce several additional negative Bethe–Hessian eigenvalues (up to three for $TS(28,22)$) and raise $\widehat{A}$, thereby generating many low‑energy directions that corrupt the spectral embedding. The resulting classification scores are Top‑1 accuracy: $0.1368\pm 3.72\times 10^{-4}$; Top‑3 accuracy: $0.2450\pm 5.63\times 10^{-5}$. The more than four‑fold drop in top‑1 performance underscores how a few detrimental trapping sets can cripple the discriminative power of graph‑based features. For reference we generated dense Erdős–Rényi graphs (no QC structure, original weight 5 and 10) with average degrees $d\approx 3$ and $d\approx 5$ (almost regular) after cycle broken. After six independent runs the observed accuracies were top-1 $0.617467\pm 3.72E-05$ and top-3 $0.8283\pm 2.54E-05$ (for weight 5) and top-1 $0.659333\pm 1.30E-05$ and top-3 $0.85396667\pm 8.4E-06$ (for weight 10). Even though these dense random graphs outperform the non‑optimised QC‑LDPC design, they are considerably less efficient (higher degree, no structured protection) and lack the flexibility offered by quasi‑cyclic constructions. Negative Bethe–Hessian modes identify substructures that generate low‑energy directions in the spectral embedding; each such direction reduces feature orthogonality. The continuous genus $\widehat{A}$ provides a quantitative measure of how “dangerous’’ a trap is: larger values correlate with more severe degradation of downstream performance. By pruning high‑genus, multi‑mode traps (e.g. $TS(4,2)$, $TS(28,22)$) we can keep the graph sparse ($\bar{w}\approx 3.6$) while retaining classification accuracy comparable to much denser Erdős–Rényi graphs, and improve top-3 accuracy important for graph ensemble (using unequal protection of features, introduce high degree columns). The presented design methodology—selecting protographs, eliminating harmful trapping sets, and fine‑tuning via Bayesian optimization—yields QC‑LDPC graphs that combine structured protection, and strong discriminative power. All auxiliary material (list of small weight TS, TS submatrix, parity-check matrix, eigenvalue trajectories, topology invariants and related quadratic form signatures, source code, etc) is available in the public repository [40].
Приложение B Trapping sets as topological defects in the spectral embedding
@@ A.6 Influence of Topological Invariants on Spectral Embedding
In this appendix, we give a completely rigorous step-by-step proof that any trapping‑set subgraph $\mathcal{T}$ of a QCLDPC Tanner graph creates a low‑energy (negative) mode of the Bethe–Hessian $H_{\beta,J}$ and, equivalently, produces a pole of the Ihara–Bass zeta function. The argument is split into four formal statements.
@@
  • $H\in\{0,1\}^{m\times n}$ – binary incidence (parity‑check) matrix of a QC‑LDPC code; rows = check nodes, columns = variable nodes.
@@
$B_{ij}^{\mathcal{T}}(\beta)$ $=\frac{\tanh(\beta J_{ij})}{1-\tanh^{2}(\beta J_{ij})},$ $i\neq j,\;(i,j)\in E_{\mathcal{T}},D^{\mathcal{T}}_{ii}(\beta)$ $=1+\sum_{j:(i,j)\in E_{\mathcal{T}}}\frac{\tanh^{2}(\beta J_{ij})}{1-\tanh^{2}(\beta J_{ij})}.$
@@
Let $C\subseteq\mathcal{T}$ be a simple cycle of length $\ell\geq 2$ satisfying (6) on each of its edges. Then the restriction $H_{\beta_{N},J}^{C}$ possesses exactly one negative eigenvalue; all other eigenvalues are non‑negative. The corresponding (up to a gauge transformation) eigenvector is $v=(1,-1,1,-1,\dots)\in\mathbb{R}^{\ell},$ i.e. it alternates sign along the cycle.
Доказательство.
Because the Tanner graph of an LDPC code is bipartite, every cycle $C$ is even; write $\ell=2r$. Hence a gauge transformation $x_{i}\;\mapsto\;s_{i}\,x_{i},\qquad s_{i}\in\{\pm 1\},$ can be chosen so that after the change of variables every edge of $C$ has positive coupling: $\tanh(\beta J_{ij})=+\sqrt{1-\varepsilon},\qquad\varepsilon:=1-\tanh^{2}(\beta J_{ij})>0.$ The parameter $\varepsilon$ depends on the edge, but for all edges of $C$ we have $\varepsilon\downarrow 0$ as $\beta\to\beta_{N}$. For an edge $(i,j)\in C$ we compute $\frac{\tanh^{2}(\beta J_{ij})}{1-\tanh^{2}(\beta J_{ij})}=\frac{1-\varepsilon}{\varepsilon},\qquad\frac{\tanh(\beta J_{ij})}{1-\tanh^{2}(\beta J_{ij})}=\frac{\sqrt{1-\varepsilon}}{\varepsilon}.$ All other edges (those not belonging to $C$) contribute bounded terms that remain finite as $\varepsilon\to 0$; we denote their total contribution by a matrix $R_{\varepsilon}$ satisfying $R_{\varepsilon}=O(1)$. Let $A_{C}$ be the adjacency matrix of the cycle (entries equal to 1 on the two neighboring positions, zero otherwise). Using Lemma B.1 we obtain for the restriction to $C$: $H_{\beta,J}^{C}=\Bigl(1+\frac{2(1-\varepsilon)}{\varepsilon}\Bigr)I-\frac{\sqrt{1-\varepsilon}}{\varepsilon}A_{C}+R_{\varepsilon}.$ Factorizing $1/\varepsilon$ and expanding $\sqrt{1-\varepsilon}=1+O(\varepsilon)$ gives the asymptotic expansion
@@
Because $\mu_{0}=0$, 8 gives $\lambda_{0}(\varepsilon)\to-1$, whereas for every $k\neq 0$ we have $\mu_{k}>0$ and thus $\lambda_{k}(\varepsilon)\to+\infty$ as $\varepsilon\downarrow 0$. Consequently, in the limit $\beta\to\beta_{N}$ there is exactly one negative eigenvalue (tending to 1-1) and all remaining eigenvalues are non‑negative. The eigenvector belonging to $\mu_{0}=0$ is the constant vector in the gauge‑ transformed basis. Undoing the gauge transformation multiplies each component by the sign factor $s_{i}$, which yields precisely the alternating pattern $v=(1,-1,1,-1,\dots)$. Since the cycle length is even, this vector is well defined. ∎
@@
Доказательство.
Decompose $\mathcal{T}$ into its $c(\mathcal{T})$ connected components $\mathcal{T}^{(1)},\dots,\mathcal{T}^{(c)}$. Because $H_{\beta_{N},J}^{\mathcal{T}}$ is block‑diagonal with respect to this decomposition, the total number of negative eigenvalues is the sum over the components. Fix a component $\mathcal{C}$. Choose a spanning forest $F$ of $\mathcal{C}$; the edges of $F$ form a tree and thus contribute no negative eigenvalue (its Bethe–Hessian reduces to a positive semidefinite matrix, see Lemma B.1). Every edge $e\in E(\mathcal{C})\setminus F$ creates exactly one independent cycle when added to $F$. By Proposition 1 each such cycle produces one negative eigenvalue of $H_{\beta_{N},J}^{\mathcal{C}}$, and different edges generate linearly independent cycles, hence distinct negative eigenvalues. Therefore, $q_{\mathcal{C}}=|E(\mathcal{C})|-|V(\mathcal{C})|+1.$ Summing over all components gives $q_{\mathcal{T}}=\sum_{c=1}^{c(\mathcal{T})}\!\bigl(|E(\mathcal{T}^{(c)})|-|V(\mathcal{T}^{(c)})|+1\bigr)=|E(\mathcal{T})|-|V(\mathcal{T})|+c(\mathcal{T})=\beta_{1}(\mathcal{T}),$ which is precisely the claimed equality. ∎
@@
We recall the non‑backtracking (Hashimoto) matrix and the Ihara–Bass zeta function.
@@
For a finite graph $\Gamma$ let $\mathcal{B}(\Gamma)$ be its non‑backtracking operator. The Ihara–Bass zeta function is $\zeta_{\Gamma}(u)=\prod_{[C]}\bigl(1-u^{\,\ell(C)}\bigr)^{-1},$ where the product runs over all primitive (non‑repeating) cycles $[C]$ and $\ell(C)$ denotes their length.
A classical identity (see e.g. [12]) links $\zeta_{\Gamma}$, the non‑backtracking matrix, and the Bethe–Hessian:
@@
Corollary B.2.1.
Let $\Gamma$ be any graph. If for some $u_{c}$ we have $\det(H_{\beta,J(u_{c})})=0$, then $u_{c}$ is a pole of the Ihara–Bass zeta function. In particular, every negative eigenvalue identified in Theorem B.2 produces a pole at $u_{c}=\tanh(\beta_{N}J_{e})$, where $J_{e}$ is the coupling on the corresponding cycle edge.
Доказательство.
Equation (9) shows that $\zeta_{\Gamma}(u)$ diverges exactly when $\det(H_{\beta,J(u)})=0$. By Theorem B.2, at the Nishimori temperature each independent cycle forces an eigenvalue of $H_{\beta_{N},J}$ to cross zero (from negative for $\beta>\beta_{N}$ to positive for $\beta<\beta_{N}$). Hence the corresponding $u_{c}=\tanh(\beta_{N}J)$ is a pole of $\zeta_{\Gamma}(u)$. ∎
The exact correspondence proved above relies on the strict Nishimori condition $\tanh^{2}(\beta_{N}J_{ij})=1$ for every edge belonging to a cycle. In practice this holds only for uniform couplings $|J_{ij}|=\infty$ or when the distribution of $J$ is chosen so that the Nishimori line coincides with $\tanh^{2}=1$. Nevertheless, the mechanism is robust: Even if $\tanh^{2}(\beta J_{ij})<1$ but close to 1 on a cycle, the two terms in (5) become large with opposite sign. The matrix $H_{\beta,J}^{C}$ then has one very small eigenvalue, which becomes negative as soon as $\beta$ exceeds a critical value that is close to $\beta_{N}$. Consequently, away from the exact Nishimori line the number of negative eigenvalues need not equal $\beta_{1}(\mathcal{T})$, but it is still bounded below by the number of balanced even cycles and above by the total number of independent cycles. In other words, $\beta_{1}(\mathcal{T})-\delta\leq q_{\mathcal{T}}\leq\beta_{1}(\mathcal{T})$ for a small defect $\delta$ that vanishes as the Nishimori condition is approached. The Ihara–Bass identity (9) holds for any graph, irrespective of the values of $J$. Therefore trapping sets always generate poles of $\zeta_{\Gamma}$; at the exact Nishimori temperature these poles correspond to eigenvalue crossings of the Bethe–Hessian. Thus the result proved in Steps 1–4 is a sharp statement for the idealized Nishimori point, while the qualitative picture (cycles $\rightarrow$ low‑energy modes $\rightarrow$ zeta‑function poles) persists under realistic, approximate conditions.